# Development of Neural Network-Based Optimal Control Pulse Generator for Quantum Logic Gates Using the GRAPE Algorithm in NMR Quantum Computer

### v1

@@ IV Testing the neural network
To evaluate the performance of the developed neural network, we employed various numerical illustrations. Firstly, We began by uniformly selecting points on the sphere (Using the cosine inverse approach) to determine the corresponding rotation axis of the unitary operations in Bloch sphere. We also uniformly selected the angle of the rotation around that axis. This process was repeated 15,000 times. Finally, these unitary matrices were fed into the trained network. The fidelity was then calculated using the standard definition:
$F(\rho,\sigma)=Tr({\rho\sigma}).$ (4)
Next, we plotted the histogram of the fidelity, as illustrated in Fig.6, which demonstrates the perfect quality of the results.
Figure 6: Histogram representing the frequency of the fidelities vs fidelity of the test results from the output of the neural network. The majority of the gates are implemented with a fidelity greater than 0.9.
In Table.1, you’ll find information on the generated pulses. As illustrated, this method is significantly faster than the standard GRAPE approach, which typically takes at least 15 seconds. Consequently, this method is approximately three orders of magnitude faster, and can be integrated as part of the compiler in quantum computer architectures.
@@
$U_{x}(\theta)=He^{-i\theta\sigma_{x}^{(1)}},U_{y}(\theta)=He^{-i\theta\sigma_{y}^{(1)}},U_{z}(\theta)=He^{-i\theta\sigma_{z}^{(1)}}.$ (5)
@@
Figure 7: Fidelity of unitary transformations given in Eq.5 vs $\theta$, as derived from the output of the trained network, where parts a, b and c correspond to operations $U_{x}(\theta)$, $U_{y}(\theta)$ and $U_{z}(\theta)$, respectively.
@@
Figure 12: Experimental results of the NMR spectrum around qubit one after applying Neural Network-generated pulses for the gates in Eq.6 on pseudo-pure state for $\theta=\frac{\pi}{4}$ and represented $\alpha$. The blue curve represents the real part of the Fourier transform, the red curve depicts the imaginary part.
Figure 13: Simulation results of the NMR spectrum around qubit one after applying gates in Eq.6 on pseudo-pure state for different values of $\alpha$. The blue curve represents the real part of the Fourier transform, the orange curve depicts the imaginary part.
The mean and standard deviation of the fidelites of implementing the unitary transformation given in Eq.6 with our network are 95% and 8%, respectively.
V Experimental Test
To ensure that the neural network will perform properly in real experiments, we experimentally implemented the generated pulses for quantum logic gates of Eq.6 on a real benchtop NMR system. In NMR spectroscopy, the average value of the magnetization in the direction prependicular to $B_{0}$, is measured and subsequently subjected to a Fourier transformation.It has been shown that within the frequency domain, the observed peaks correspond proportionally to specific elements of the system’s density matrix [45, 2, 3]. Our NMR system is a $1Tesla$ benchtop spectrometer with $C_{2}F_{3}I$ molecule, with fluorine nuclei serving as qubits. At $1Tesla$, we observed the following constants of the experiment:
$\omega_{1}-\omega_{rf}=-1375\,\text{Hz},$ (7)
$\omega_{2}-\omega_{rf}=56\,\text{Hz},$
$\omega_{3}-\omega_{rf}=-1035\,\text{Hz}$
$\omega_{rf}=40\,\text{MHz},$ (8)
$J_{1,2}=-67\,\text{Hz},$
$J_{1,3}=28\,\text{Hz},$
$J_{2,3}=38\,\text{Hz}$
where the term $\omega_{i}$ represents the Larmor frequency for the $i^{th}$ nucleus. In Fig.10, the output spectrum of the simple square shape pulse applied to the thermal state is depicted. In the context of NMR spectroscopy for a 3-qubit system with fully coupled nuclei, one would expect to observe 12 distinct peaks [45, 2, 3]. This phenomenon is clearly demonstrated in the output spectrum. To demonstrate the precision of the neural network, we generated pulses by using various values of $\alpha$ and $\theta=\frac{\pi}{4}$ for the operations of the form Eq.6. Fig.13 represents the simulation results of the spectrum around qubit-1 after applying gates presented in Eq.6 for different $\alpha$ values on pseudo-pure state. Fig.13 illustrates the theoretical behavior of both real and imaginary components of the NMR spectrum. In Fig.12 the experimental result for these pulses are presented. Comparing the experimental results with the simulations for various values, one can observe the consistency of the phase behavior between the real and imaginary parts. Now, we perform integration on the spectrum of real and imaginary functions in Fig.12 and Fig.13, and subsequently define the corresponding phase of the output signal as follows:
@@ V Experimental Test
$Phase=\tan^{-1}(\frac{Im}{Re}),$ (10)
where $f_{1}$ to $f_{2}$ represents the frequency range in which the Larmor frequency for the desired spin exclusively falls within this range—excluding other spins—as we aim to calculate the intensity related to this spin, and ‘im’ and ‘re’ correspond to the imaginary and real parts of the spectrum in Fig.12 and 13. By calculating the Phase in Eq.10, for different values of $\alpha$ in Fig.12 and 13, we can obtain the results shown in Fig.14. This similarity in the simulation and experimental outcomes suggests that the model is sufficiently precise at this scale, and also the Neural Network Generated-pulses perform admirably.
@@
For the theoretical discussion, consider applying the gate in Eq.6 to the state |$\ket{000}$:
$e^{-i\frac{\pi}{4}(\cos(\alpha)\sigma_{x}^{(1)}+\sin(\alpha)\sigma_{y}^{(1)})}\ket{000}=$ (11)
$\frac{1}{\sqrt{2}}(\ket{000}+(-i\cos(\alpha)+\sin(\alpha))\ket{100})$
By calculating the output signal of NMR, we observe that the integral of the output Fourier transform of the signal for the aforementioned state (For an equivalent state relative to the pseudo-pure state.), around the frequency related to qubit one is [45, 2, 3]:
$Im=-\cos(\alpha),Re=\sin(\alpha)\rightarrow$ (12)
$Phase=\tan^{-1}(-\cot(\alpha))$
Now, it is clear why there is a linear relationship in Fig. 14 between the defined phase in Eq.10 and $\alpha$ in Eq. 6.
@@ VI Conclusion
In summary, we demonstrated the feasibility of extending optimal control theory for implementing arbitrary quantum logic gates through neural networks and illustrated the employment of the GRAPE algorithm to train the network. We demonstrated that using this approach, the cosine similarity between GRAPE-generated pulses used as training samples increases when the GRAPE algorithm is initiated from the same point, facilitating the learning process. After implementing the network, we tested it using numerical simulations, thereby verifying the success of the method. The feasibility of directly implementing arbitrary single-qubit gates without decomposing them into basic gates, helps achieve logical gates with higher fidelity and fewer steps. This approach significantly enhances circuit depth, which is particularly crucial in the NISQ era. When combined with the GRAPE-generated pulse for the CNOT gate, it creates an effective set of universal gates capable of implementing more complex algorithms in today’s quantum computers. Moreover, the approximate real-time generation of control pulses for arbitrary operations allows the network to be integrated into the compiler within quantum computer architecture. Finally, we experimentally validated the effectiveness of the output control pulses generated by the developed neural network. For future research, it is important to expand this concept to arbitrary two-qubit gates (i.e., Controlled-U gates) and explore the feasibility of designing a neural network for a larger number of qubits. This is crucial because, as the number of qubits increases, the system’s dimensionality grows exponentially. This growth could hinder our ability to compute such large-dimensional matrices, rendering the calculation of overall dynamics infeasible.

### v2

@@ IV Testing the neural network
To evaluate the performance of the developed neural network, we conduct a series of numerical experiments. For the primary assessment, the network is tested on a dataset of 15,000 uniformly distributed $2\times 2$ unitary matrices generated using Algorithm 2.
Finally, we feed these unitary matrices into the trained network and compute the operation that would be applied to the system upon implementation of the neural network-generated pulse. To determine the total resultant operation, we employ the same approach used in the GRAPE algorithm [43], as detailed in Eqs.5 and 6. The fidelity was then calculated between $U_{Target}$, the unitary operation intended as the input to the neural network, and $V$, the unitary operation actually implemented by the pulse generated by the neural network, using the definition:
$F(U_{Target},V)=Tr(U_{Target}^{\dagger}V).$ (7)
Next, we plot the histogram of the fidelity, as illustrated in Fig.6, which demonstrates the perfect quality of the results.
Figure 6: Histogram illustrating the distribution of fidelities in the test results obtained from the neural network’s output. The data-set comprises 15,000 samples, with the majority of gates achieving a fidelity greater than 0.9.
In Table.1, you’$l$ find information on the generated pulses. As illustrated, this method is significantly faster than the standard GRAPE approach, which typically takes at least 15 seconds. Consequently, this method is approximately three orders of magnitude faster, and can be integrated as part of the compiler in quantum computer architectures.
@@
$\begin{split}U_{x}(\theta)&=He^{-i\theta\sigma_{x}^{(1)}},\\ U_{y}(\theta)&=He^{-i\theta\sigma_{y}^{(1)}},\\ U_{z}(\theta)&=He^{-i\theta\sigma_{z}^{(1)}}.\end{split}$ (8)
@@
Figure 7: Fidelity of unitary transformations given in Eq. 8 as a function of $\theta$, derived from the output of the trained network. Panels (a), (b), and (c) correspond to the operations $U_{x}(\theta)$, $U_{y}(\theta)$, and $U_{z}(\theta)$, respectively.
@@
The mean and standard deviation of the fidelities for implementing the unitary transformations described in Eq. (6) using our network are $95\%$ and $8\%$, respectively. As illustrated in Figs. 7 and 8, noticeable dips in fidelity are observed. A possible explanation is that limited sampling coverage, combined with the inherent nonlinearity of the underlying physical system, may be more pronounced for many of these low-fidelity gates, thereby making the training process more challenging. Nonetheless, increasing the network depth did not lead to a significant improvement in fidelity, suggesting that the learning dynamics have saturated under the current approach.
V Experimental Test
To ensure that the neural network will perform properly in real experiments, we experimentally implemented the generated pulses for quantum logic gates of Eq.9 on a real benchtop NMR system. In NMR spectroscopy, the average value of the magnetization in the plan perpendicular to $B_{0}$, is measured and subsequently subjected to a Fourier transformation. It has been shown that within the frequency domain, the observed peaks correspond proportionally to specific elements of the system’s density matrix [2, 3]. This process is thoroughly discussed in quantum state tomography on an NMR systems [46]. Our NMR system is a 11 TeslaTesla benchtop spectrometer with $C_{2}F_{3}I$ molecule, with fluorine nuclei serving as qubits. At 11 TeslaTesla, we observed the following constants of the experiment:
[…]
@@ V Experimental Test
where $\omega_{i}$ represents the Larmor frequency of qubit $i$, $\omega_{rf}$ is the frequency of applied RF field and $J_{ij}$ represents the J-coupling coefficients between qubit $i$ and qubit $j$. Fig.10 depicts the output spectrum of the system under applying the simple square shape pulse to the thermal state of the sample. For wide spectral width of spins one should check that square shaped RF pulse must able to homogeneously excite all the three spins. one example of uniform excitation of spins with wide spectral width can be found in [47]. In the context of NMR spectroscopy for a 3-qubit system with fully coupled nuclei, one would expect to observe 12 distinct peaks [46, 2, 3] which is clearly demonstrated in the output spectrum of Fig.10. To demonstrate the precision of the neural network, we generate pulses corresponding to gates of the form given in Eq.9 with various values of $\alpha$ and $\theta=\frac{\pi}{4}$ and then consider the spectrum around qubit 1 after applying the generated pulses on the pseudo-pure state (PPS) [48].
@@
It has been shown that the integrals of the real and imaginary components of the output spectrum near the resonance frequencies determine the density matrix elements of the system [2, 3]. Figure 12 presents the simulated real and imaginary parts of the NMR output spectrum for a pseudo-pure state associated with qubit one, under applied pulses corresponding to the operations defined in Eq. 9, with $\theta=\frac{\pi}{4}$ and various values of $\alpha$. Since the single-qubit gate is applied exclusively to qubit one, we focus on monitoring the spectral evolution near frequencies close to its Larmor frequency. Figure 13 shows the corresponding experimental results.
Figure 12: Simulation results of the NMR spectrum around qubit one after applying gates in Eq.9 on pseudo-pure state for $\theta=\frac{\pi}{4}$ and different values of $\alpha$. The blue curve represents the real part of the Fourier transform, the orange curve depicts the imaginary part.
Figure 13: Experimental results of the NMR spectrum around frequencies close to the Larmor frequency of the qubit one, after applying Neural Network-generated pulses for the gates in Eq.9 on pseudo-pure state for $\theta=\frac{\pi}{4}$ and represented $\alpha$. The blue curve represents the real part of the Fourier transform, the red curve depicts the imaginary part of the spectrum.
It is naturally expected that the experimental spectrum deviates from the theoretical one, as the theoretical model assumes a closed quantum system. In reality, however, the system is not closed. As a result, the pulse shapes derived from the theoretical model do not perform perfectly in practical scenarios, leading to discrepancies between the theoretical and experimental spectra under identical operations. It is important to emphasize that quantum control and pulse shaping are employed to enhance gate fidelities by controlling unwanted interactions within closed quantum systems. However, these techniques do not address errors arising from environmental interactions and external sources of decoherence in open quantum systems. Several environmental factors contribute to the deviations observed between the predicted and measured spectra. These include non-homogeneous static magnetic fields, temporal fluctuations in magnetic field strength due to temperature variations, stray electromagnetic fields, and phase or amplitude noise in the RF signals. For such systems, solutions like quantum error correction and error mitigation are required. As discussed in the introduction, in closed quantum systems—particularly in NMR systems—the dominant sources of error include unwanted interactions of RF pulses with non-target qubits (especially in weak-field benchtop NMR systems) and unavoidable two-qubit interactions during single-qubit gate operations. In these cases, pulse shapes are engineered so that the target qubit undergoes the intended operation, while non-target qubits also evolve but return to their initial states by the end of the gate operation. In other words, the non-target qubits effectively experience an identity operation.
[…]
Next, we consider the NMR signal spectrum corresponding to the above state and perform integration over its real and imaginary components within the frequency range associated with qubit 1:
$\text{Im}=\int_{f_{1}}^{f_{2}}\text{im}(f)\,df,\quad\text{Re}=\int_{f_{1}}^{f_{2}}\text{re}(f)\,df,$ (13)
where $f_{1}$ to $f_{2}$ denotes the frequency range that exclusively contains the Larmor frequency of the target spin (qubit 1), excluding contributions from other spins (qubits 2 and 3). This selective integration allows us to isolate and quantify the signal intensity associated with qubit 1. The functions $\text{im}(f)$ and $\text{re}(f)$ represent the imaginary and real parts of the NMR spectrum, respectively.
We then define the phase of the output signal as
$\text{Phase}=\tan^{-1}\left(\frac{\text{Im}}{\text{Re}}\right).$ (14)
By calculating the output NMR signal and its Fourier transform, as detailed in Appendix A, we find that the quantities ImIm, ReRe and PhasePhase for the state given in Eq. 12, evaluated around the frequency corresponding to qubit 1, yield the following results:
$\text{Im}=-\cos(\alpha),\quad\text{Re}=\sin(\alpha)$ (15)
$\quad\Rightarrow\quad\text{Phase}=\tan^{-1}(-\cot(\alpha)),$
which demonstrates a linear relationship between the defined phase in Eq. 14 and the parameter $\alpha$ in Eq. 12.
Finally, by obtaining the experimental NMR output spectrum corresponding to the state defined in Eq. 12 for various values of $\alpha$ we compute the phase as described in Eq. 14 for each value of $\alpha$. The results are presented in Fig.14, alongside the theoretical predictions obtained from simulations. The close agreement between the experimental data and simulated results indicates that the model exhibits a good degree of accuracy at this scale, and that the neural network–generated pulses demonstrate reliable performance.
@@ VI Conclusion
In summary, we demonstrated the feasibility of extending optimal control theory for implementing arbitrary quantum logic gates through neural networks and illustrated the employment of the GRAPE algorithm to train the network. We demonstrated that using this approach, the cosine similarity between GRAPE-generated pulses used as training samples increases when the GRAPE algorithm is initiated from the same point, facilitating the learning process. After implementing the network, we tested it using numerical simulations, thereby verifying the success of the method. The feasibility of directly implementing arbitrary single-qubit gates without decomposing them into basic gates, helps achieve logical gates with higher fidelity and fewer steps. This approach significantly enhances circuit depth, which is particularly crucial in the NISQ era. When combined with the GRAPE-generated pulse for the CNOT gate, it creates an effective set of universal gates capable of implementing more complex algorithms in today’s quantum computers. Moreover, the approximate real-time generation of control pulses for arbitrary operations allows the network to be integrated into the compiler within quantum computer architecture. Finally, we experimentally validated the effectiveness of the output control pulses generated by the developed neural network. For future research, further investigation can be carried out to understand the nature of the dips presented in Figs.8 and 9. Utilizing different architectures, data generation procedures, and physics-informed losses could potentially lead to higher fidelities and mitigate the network’s bias toward gates that were not sufficiently learned. It is also important to expand this concept to arbitrary two-qubit gates (i.e., Controlled-U gates) and explore the feasibility of designing a neural network for a larger number of qubits. This is crucial because, as the number of qubits increases, the system’s dimensionality grows exponentially. This growth could hinder our ability to compute such large-dimensional matrices, rendering the calculation of overall dynamics infeasible.
Appendix A Derivation of Eq.15
To derive Eq.15, we provide a more detailed discussion of the NMR spectrum. The output NMR signal is proportional to the magnetization in the plane perpendicular to the static magnetic field $B_{0}$ [46]:
$\braket{M_{x}}=Tr(\rho(t)\sum^{n}_{k=1}\sigma^{(k)}_{x})$, (16)
where $\rho(t)$ is the density matrix of the system at time $t$, during the free evolution of the system governed by the Hamiltonian $H_{free}$ in Eq.2. Using this, we obtain the following expression for $\braket{M_{x}}$:
$Tr(e^{\frac{-it}{\hbar}H_{free}}\rho_{0}e^{\frac{it}{\hbar}H_{free}}\sum^{n}_{k=1}\sigma^{(k)}_{x})$. (17)
Expanding $\rho_{0}$, the initial state of the system, in the eigenstates of $H_{free}$, we obtain:
[…]
where $E_{l}$ values are eigenvalues of $H_{free}$ and when we expanded $\rho_{0}$, we only considered the traceless part in Eq.12, because the other part plays no role in the spectrum. Consequently, by substituting Eq.12 in the Eq.17, we obtain the magnetization signal as:
$\frac{1}{2}\left(-ie^{i\alpha}e^{\frac{-it}{\hbar}(E_{000}-E_{100})}+ie^{-i\alpha}e^{\frac{it}{\hbar}(E_{000}-E_{100})}\right)=$ (19)
$\mathrm{Re}\left(-ie^{i\alpha}e^{-i\omega t}\right)=\sin(\alpha)\cos(\omega t)-\cos(\alpha)\sin(\omega t),$
where,
$\omega=\frac{1}{\hbar}(E_{000}-E_{100}),$ (20)
By considering the resultant Fourier transform of the above signal, it is evident that the real part of the transform equals $\sin(\alpha)\delta(v-\omega)$, while the imaginary part equals -$\cos(\alpha)\delta(v-\omega)$. This approach allows us to demonstrate the validity of Eq.15. Under nonideal conditions and in real experiments, some form of damping occurs in Eq.17 [2]. This means that we can estimate the final signal as:
$e^{-t/T}(\sin(\alpha)\cos(\omega t)-\cos(\alpha)\sin(\omega t)).$ (21)
Again, by applying the Fourier transform to Eq.21, along with the detailed discussion in quantum state tomography on an NMR system [51, 46, 2, 3], we can demonstrate that the terms defined in Eq.17 justify Eq.15. By integrating the transformed signal, we obtain results that closely resemble the ideal situation.